solve each quadratic in form equation

5. 2(x^2-5)^2-13(x^2-5)+20=0
1
Expert's answer
2012-12-10T11:06:59-0500
let x^2-5=t then
2t^2-13t+20=0
D=13^2-4*2*20=9
t=(13+-3)/4
t=4 or t=5/2x^2=t+5
so
x^2=9 or x^2=15/2
x1=-3, x2=3, x3=-sqrt(15/2), x4=sqrt(15/2)